# Exploring the Robustness of Language Models for Tabular Question Answering via Attention Analysis

### v3

@@ 4 Attention Analysis
Figure 2: Heatmap showing Spearman correlation between changes in attention entropy and EM difference across all attention heads and layers in the LLama-8B-Instruct model on the WTQ dataset.
@@ 4.1 Effect of Perturbations on Attention Maps
We examine how varying severities of structural perturbations influence attention weights using the Llama3-8B-Instruct model on the WTQ dataset. For each attention head across all layers, we measure the change in attention entropy (detailed in Appendix H) between the original and perturbed table inputs. In parallel, we compute the corresponding change in Exact Match (EM) scores, capturing the performance impact of each perturbation.
@@
4.2 Layer-Wise Analysis of Attention Head Sensitivity to Perturbations
Figure 3: Layer-wise Spearman correlation between aggregated attention entropy change (averaged over heads) and EM difference.
@@ 4.2 Layer-Wise Analysis of Attention Head Sensitivity to Perturbations
While averaging attention metrics across heads provides a general understanding of model behavior, prior research emphasizes that individual attention heads and specific layers, particularly the middle layers, significantly contribute to encoding task-relevant information Barbero et al. (2025); Zhao et al. (2024). To investigate this at a finer granularity, we perform a layer-wise analysis of how perturbation-induced changes in attention entropy correlate with EM score degradation.
@@
Specifically, we use the Llama3-8B-Instruct model on the WTQ dataset to compute the Spearman correlation between entropy changes (original vs. perturbed inputs) and corresponding EM differences for each attention head across all layers. Figure 2 demonstrates that correlations peak predominantly in the middle layers and remain consistently elevated through these central layers. This indicates that perturbation-induced shifts in attention distribution within middle layers are strongly predictive of performance degradation. Additional experiments are included within the Appendix H.
To further substantiate this pattern, we aggregate entropy changes across all heads within each layer and analyze the correlation with EM differences (Figure 3). This aggregated analysis reinforces the observation of heightened sensitivity in mid-layer attention mechanisms.Collectively, these results highlight that perturbation sensitivity is not uniformly distributed across model layers. Instead, the middle layers, which typically engage in critical representational refinement, are particularly vulnerable to attention instability and associated performance degradation.
@@ 5 Conclusion
Our study offers a comprehensive robustness analysis of LLMs under various perturbations for TQA. While larger, instruction-tuned models show improved performance, they remain highly sensitive to structural and value-based disruption. We also uncover domain biases, where models perform well on WTQ without tables yet struggle on more specialized datasets like TAT-QA and SCITAB. We show that shifts in attention entropy, particularly in middle layers, correlate with performance degradation. This work highlights the importance of interpretability and robustness of LLMs for TQA.
@@ 6 Limitation
Although we provide extensive evaluation of LLMs on WTQ, TAT-QA, and SCITAB datasets, it is possible to include a broader range of datasets for a more comprehensive comparison that would highlight the generalizability of our method for both domain-specific datasets and Wikipedia-based datasets. While we anticipate that similar performance could be achieved with other tasks, such as table summarization, future work should include extensive analysis across various tasks and datasets to validate the assumption. Moreover, our study did not involve any structurally aware or fine-tuned models for tabular datasets. It is plausible that fine-tuning and structurally enhanced models could significantly impact the performance of different models. Additionally, our evaluation relied on exact match accuracy for assessing the text generation model’s performance. This metric, while useful, limits the scope of evaluation for the question answering task. Future studies should employ more nuanced evaluation metrics to better assess the robustness of the models in TQA tasks. Moreover, we only conduct a case study of attention analysis with small models because of computation cost.
Appendix A Example Prompt
@@
Figure 5: Example of a prompt with answer for WTQ dataset for Data Type Preserving Perturbation. In comparison to Fig 4, we replace correct answer(Bankok, Thailand) with a fake answer(Beijing).
Figure 6: Example of a prompt with answer for WTQ dataset for Random Value Perturbation. Here, we replace correct answer(Bankok, Thailand) with an abstract random value (r@nD0m v@1u3).
Figure 7: Example of a prompt with answer for WTQ dataset for Random Value Perturbation. We remove the correct answer(Bankok, Thailand).
Appendix B Generating Data Type Preserving Perturbation Dataset
@@
Appendix C Evaluation Dataset Size
We select three different datasets for comparison, WTQ, TAT-QA, and SCITAB dataset, with different numbers of evaluation datasets as described in Table 2. For fair comparison, we limit the number of cell elements($<150$) within the table for both datasets. Similarly, for Value Perturbation, some queries relate to over all structure of the table, hence we filter only those tables that contains the answer value for the given query.
@@ Appendix C Evaluation Dataset Size
Appendix D Models
We selected recent open-source models that have been extensively studied. Table 3 lists all the models we considered with their parameter size and their date of release.
@@ Appendix D Models
Large Model(Larger than 40B parameter)
@@
Appendix E Evaluation Metrics
We employ three primary metrics to evaluate model performance under various perturbations in Tabular Question Answering (TQA) tasks.
Exact Match Accuracy (EM): This metric calculates the proportion of instances for which the predicted answer exactly matches the ground truth answer. Formally, if $N$ is the total number of instances, and $\text{correct}(i)$ is an indicator function that is $1$ if the prediction for instance $i$ is correct and 00 otherwise, then:
$\text{EM}=\frac{\sum_{i=1}^{N}\text{correct}(i)}{N}.$
Exact Match Difference (Emd): Let $\text{EM}_{\text{orig}}$ be the EM on the original (unperturbed) dataset, and $\text{EM}_{\text{perturbed}}$ be the EM after applying a perturbation. The Emd quantifies the change in EM due to perturbations:
$\text{Emd}=\text{EM}_{\text{perturbed}}-\text{EM}_{\text{orig}}.$
Variation Percentage (VP): This metric measures to what extent predictions change after applying perturbations. Given that, C2W is the count of correct before perturbation and wrong after, and W2C is the count of wrong before perturbation and correct after. Given $N$ as the total number of instances, the variation percentage is:
$\text{VP}=\frac{C2W+W2C}{N}.$
@@ Appendix F Performance over Table Complexity
Figure 9: The average EM score for the different structural perturbation over the different size of tables on WTQ, TAT-QA and SCITAB datasets
Figure 10: The average EM scores of the original models and fine-tuned for instruction or conversation models across various table augmentation techniques for WTQ, TAT-QA and SCITAB dataset
@@
Across all three datasets, these results highlight that large language models, though capable, display uneven robustness to structural manipulations of tables. Although instruction tuning and larger model scales improve performance, structural changes continue to pose challenges, indicating the need for more robust, structure-aware approaches for reliable table comprehension.

### v4

@@ 5 Attention Analysis
We utilize attention entropy as a quantitative measure of how attention distributions respond to structural perturbations. Attention entropy reflects the dispersion of attention weights across input tokens, serving as an indicator of how focused or diffuse the model’s attention becomes under disruption. In the context of tabular data, where semantic understanding hinges on precise structural alignment, operations like row and column swaps inherently distort positional cues. These distortions lead to measurable shifts in entropy, making it a particularly effective diagnostic for attention misalignment. By tracking how entropy changes in response to perturbations, we capture the extent to which structural instability propagates through the attention mechanism. This perspective adds on prior work showing the role of entropy in capturing context encoding diversity Zhang et al. (2025), stabilizing attention dynamics during training Zhai et al. (2023), and improving generalization under limited data conditions Araabi et al. (2024), and extends these insights to structured, table-based inputs.
@@ 5.1 Effect of Perturbations on Attention Maps
We examine how varying severities of structural perturbations influence attention weights using the Llama3-8B-Instruct model on the WTQ dataset. For each attention head across all layers, we measure the change in attention entropy between the original and perturbed table inputs. In parallel, we compute the corresponding change in Exact Match (EM) scores, capturing the performance impact of each perturbation.
@@
5.2 Analysis of Individual Attention Head Sensitivity to Perturbations
Figure 5: Heatmap showing Spearman correlation between changes in attention entropy and EM difference across all attention heads and layers in the Llama3-8B-Instruct model on the a) WTQ, b) TAT-QA, and c) SCITAB dataset
@@ 5.2 Analysis of Individual Attention Head Sensitivity to Perturbations
While averaging attention metrics across heads provides a general understanding of model behavior, prior research emphasizes that individual attention heads and specific layers, particularly the middle layers, significantly contribute to encoding task-relevant information (Barbero et al., 2025; Zhao et al., 2024). To investigate this at a finer granularity, we perform a layer-wise analysis of how perturbation-induced changes in attention entropy correlate with EM score degradation.
@@
Specifically, we use the Llama3-8B-Instruct model on the WTQ dataset to compute the Spearman correlation between entropy changes (original vs. perturbed inputs) and corresponding EM differences for each attention head across all layers. Figure 5(a) demonstrates that correlations peak predominantly in the middle layers and remain consistently elevated through these central layers. This indicates that perturbation-induced shifts in attention distribution within middle layers are strongly predictive of performance degradation. The sharp correlation peak in WTQ suggests that the model relies heavily on mid-layer representations to align tabular structures with natural language queries, particularly when interpreting entity references and table schema. Distinct, though narrower, spikes at the input layer (0) and the output layers (30303131) further reveal that both the earliest token-encoding stage and the final representational consolidation phase are also vulnerable to structural perturbations.
Figure 5(b) and 5(c) extend the analysis to TAT-QA and SCITAB, respectively. The results reveal distinct patterns tied to domain characteristics. For TAT-QA, correlations are elevated not only in the middle layers but also extend into the upper layers of the model. In contrast, SCITAB exhibits a more sharply localized pattern, with peak correlations tightly concentrated within the middle layers. Although we observe a high correlation in these middle layers, the contrastive negative correlations are primarily due to the significantly poor EM performance on the SCITAB dataset. These findings reinforce the centrality of middle-layer attention mechanisms in tabular reasoning tasks, while also highlighting domain-specific variations in the vertical distribution of sensitivity. Additional experiments are included within the Appendix H for other models(Llama3-8B, Mistral-7B, and Mistral-7B-Instruct).
5.3 Correlation within Individual Attention
Figure 6: (a) Heatmap showing Spearman correlation between changes in attention entropy and EM difference across all attention heads and layers in the Llama-8B-Instruct model on the WTQ dataset, unsorted to depict actual correlation over each attention head. (b-c) Scatter plot showing the correlation between changes in attention entropy and EM scores across five structural perturbations for the WTQ dataset. (b) The plot highlights a strong positive correlation for Layer 0, Head 10, indicating high sensitivity of this head to structural perturbations. (c) In contrast, Layer 19, Head 23 shows the least correlation. Note that each point in plots b and c averages only across data points for each specific head-layer pair.
We find that structural perturbations in tabular inputs affect the internal attention dynamics of LLMs, specifically the Llama3-8B-Instruct model on WTQ, and that these dynamics are differentially expressed across the attention heads and layers. Figure 6(a) presents an unsorted Spearman correlation heatmap between per-head changes in attention entropy and the corresponding EM score differences, revealing substantial heterogeneity across all 32 layers and 32 heads. The heatmap captures a wide range of correlation strengths, from near-zero to values approaching $\rho=0.9$, illustrating that not all attention mechanisms contribute equally to robustness. In particular, the concentration of high-correlation values in lower layers suggests that early-stage attention heads may play a foundational role in establishing reliable structural interpretations.
To unpack this variability, Figures 6(b) and 6(c) zoom in on two extreme heads identified from the heatmap. Layer 0, Head $10$ (Figure 6(b)) exhibits a strong positive relationship: higher shifts in attention entropy due to structural perturbations reliably predict larger drops in EM scores. This implies that this head is critical for encoding spatial consistency or row-column alignment in tabular inputs, and disruptions to this alignment directly degrade model accuracy. The correlation is visually evident as a tight clustering of points around a positively sloped trend line across all five structural perturbation types (Row Swap, Column Swap, Transpose, Transpose Row Swap, and Transpose Column Swap).
In contrast, Layer $19$, Head $23$ (Figure 6(c)) demonstrates minimal correlation between entropy change and EM variation. This indicates a form of functional redundancy or robustness in this head’s role; it either performs a task unrelated to structural parsing or maintains stable attention regardless of structural distortions. This wide range in sensitivity reinforces the notion that attention heads are functionally diverse and that only a subset contributes significantly to robustness under tabular perturbation.
Altogether, these findings suggest the possibility of identifying and selectively reinforcing robustness-critical attention heads through architectural tuning or fine-tuning objectives. By mapping correlation strengths across the entire model, we can isolate those mechanisms most affected by structural shifts and potentially develop strategies, such as attention regularization or selective re-weighting, that mitigate their susceptibility to disruption. This also opens avenues for pruning or interpretability studies focused on attention heads with minimal impact, offering insights into model compression or simplification without significant performance loss.
5.4 Cross-Model and Dataset Analysis of Perturbation Sensitivity
Figure 7: Spearman correlation between perturbation-induced attention-entropy change and EM degradation across 30 model–dataset pairs.
To evaluate the generality of our findings beyond the Llama3-8B-Instruct model on WTQ, we extend our correlation analysis between attention entropy change and EM performance degradation across a diverse set of model-dataset combinations. These include all the small models defined in 3, for which each models are tested on the WTQ, TAT-QA, and SCITAB datasets. For each configuration, we compute the Spearman correlation between the perturbation-induced changes in attention entropy and the corresponding drop in EM scores.
As shown in Figure 7, the positive correlation trend persists across nearly all settings. In particular, Instruct and chat variants of the models consistently show the strongest correlation values across datasets, reinforcing their heightened sensitivity to attention dispersion caused by structural perturbations. Notably, even non-instruct variants like Mistral-7B exhibit moderate to strong correlations, although the magnitude tends to be lower and more variable. We also find that the SCITAB dataset exhibits the greatest variation in correlation values, specifically due to the particularly poor performance of the Llama2-7B model on this dataset, where the baseline EM with the original table is already low enough to substantially affect the correlation analysis.
These consistent patterns across models and datasets confirm the robustness of our main claim: the severity of perturbation that induces greater shifts in attention entropy is reliably associated with a decline in model performance. This suggests that the change in attention entropy serves as a broadly applicable proxy for evaluating robustness in table-based QA models.
@@ 6 Conclusion
Our study provides a comprehensive and robust analysis of LLMs under various perturbations for TQA. While larger, instruction-tuned models show improved performance, they remain highly sensitive to structural and value-based disruption. These disruptions notably manifest as perturbations such as random row swaps, column swaps, and transpositions, highlighting vulnerabilities in their structural reasoning capabilities. We also uncover domain biases, where models perform well on WTQ without tables but struggle on more specialized datasets, such as TAT-QA and SCITAB. Specifically, the performance on WTQ even in the absence of tables indicates reliance on memorized textual patterns from pretraining, rather than genuine tabular reasoning. In contrast, specialized domains such as financial and scientific tables pose more significant challenges due to their complexity and unique domain specificity.
We demonstrate that shifts in attention entropy, particularly in middle layers, are correlated with performance degradation. This observation was supported by detailed attention-level analyses, which revealed that attention heads in middle layers are particularly critical in encoding structural information, and their instability directly contributes to errors in comprehension. Furthermore, layer-wise attention analysis reveals that certain attention heads exhibit greater sensitivity, suggesting that targeted improvements in these areas could enhance robustness.
Our findings underline the critical need for improving both the interpretability and robustness of LLMs in TQA. Building on these insights, future research should focus on developing fine-grained attention-interpretability methods and domain-adaptive processing strategies to enhance transparency, cross-domain generalization, and reliability in practical applications.
@@ 7 Limitation
Although we provide extensive evaluation of LLMs on WTQ, TAT-QA, and SCITAB datasets, it is possible to include a broader range of datasets for a more comprehensive comparison that would highlight the generalizability of our method for both domain-specific datasets and Wikipedia-based datasets. While we anticipate that similar performance could be achieved with other tasks, such as table summarization, future work should include extensive analysis across various tasks and datasets to validate the assumption. It would also be insightful to expand the analysis to make a comparison with long-context QA. Moreover, our study did not involve any detailed analysis on structurally aware or fine-tuned models for tabular datasets. It is plausible that fine-tuning and structurally enhanced models could significantly impact the performance of different models. Additionally, our evaluation relied on exact match accuracy for assessing the text generation model’s performance. This metric, while useful, limits the scope of evaluation for the question answering task. Future studies should employ more nuanced evaluation metrics to better assess the robustness of the models in TQA tasks. Moreover, we only conduct a case study of attention analysis with small models because of the computation cost.
Appendix A Dataset
We study three tabular QA datasets chosen for complementary domains and reasoning demands. WTQ (Wikipedia tables) emphasizes broad-coverage knowledge with mixed textual–numeric cells and multi-step, compositional queries over semi-structured schemas. TAT-QA (financial reports) combines dense numeric content with accompanying textual descriptions, stressing arithmetic reasoning and text–table fusion. SCITAB (scientific tables) features claims supported by measurements and categorical attributes, prioritizing structured logical inference under domain-specific vocabulary. To limit length-induced degradation and focus the analysis on table comprehension, we cap all tables at <150<150 cells (Appendix D). We further stratify results by table-size bins (Appendix F; Fig. 13) to examine how robustness and attention behavior evolve with increasing structural complexity.
WTQ dataset
WTQ features tables taken from Wikipedia and paired with natural language questions that test a model’s ability to reason across semi-structured data. This dataset is especially useful for evaluating two key aspects of question answering: the breadth of domain knowledge and the depth of logical reasoning. Unlike datasets that focus on only one of these, WTQ challenges models on both fronts. Because the tables come from a wide range of topics on Wikipedia, WTQ introduces semantic variability, which tests a model’s ability to generalize beyond surface-level patterns.
TAT-QA dataset
TAT-QA blends structured financial tables with associated unstructured textual explanations, challenging models to align table content with surrounding text. The questions frequently require arithmetic computations (e.g., difference, sum, ratio) and integration of multiple modalities. This dataset simulates real-world information extraction tasks where decision-making depends on both tabular facts and contextual narrative, making it ideal for evaluating text-table fusion capabilities and numerical robustness. Errors in TAT-QA tend to reveal how well models can track quantitative facts across modalities.
SCITAB dataset
SCITAB introduces TQA grounded in scientific inquiry, where tables often encode structured results from controlled experiments or observational studies. The dataset features complex relational structures, such as columnar trends, controlled variables, and statistical comparisons. Unlike WTQ or TAT-QA, SCITAB emphasizes scientific reasoning, requiring models to interpret claims, draw evidence-backed conclusions, and perform multi-hop inferences within a structured format. The nuanced scientific language and evidence-based justification make SCITAB a strong benchmark for testing logical rigor and attention to detail in LLMs.
Appendix B Example Prompt
@@
Figure 9: Example of a prompt with answer for WTQ dataset for Data Type Preserving Perturbation. In comparison to Fig. 8, we replace the correct answer(Bangkok, Thailand) with a fake answer(Beijing).
Figure 10: Example of a prompt with answer for WTQ dataset for Random Value Perturbation. Here, we replace the correct answer(Bangkok, Thailand) with an abstract random value (r@nD0m v@1u3).
Figure 11: Example of a prompt with answer for WTQ dataset for Random Value Perturbation. We remove the correct answer(Bangkok, Thailand).
Appendix C Generating Data Type Preserving Perturbation Dataset
@@
Appendix D Evaluation Dataset Size
@@
We select three different datasets for comparison, WTQ, TAT-QA, and SCITAB datasets, with different numbers of evaluation datasets as described in Table 2. Each dataset provides distinct characteristics, WTQ is composed of Wikipedia tables, TAT-QA centers on financial reports, and SCITAB focuses on scientific claims, which together enable a comprehensive and robust assessment. For fair comparison, we limit the number of cell elements (<150<150) within the table for both datasets. This constraint ensures that models are not disproportionately affected by excessively large tables, which could skew performance due to context window limitations. Moreover, keeping table size bounded allows for more consistent measurement of the effects of perturbations across datasets.
Similarly, for Value Perturbation, some queries relate to the overall structure of the table. Hence, we filter only those tables that contain the answer value for the given query. This filtering ensures semantic alignment between the question and the modified table, avoiding misleading evaluations where no correct answer exists in the perturbed version. In the case of the WTQ dataset, for instance, not all table-question pairs are amenable to value alterations, particularly when the table structure is insufficiently informative or lacks direct answer candidates. Altogether, this preprocessing pipeline yields a curated benchmark that is well-suited for analyzing perturbation sensitivity while controlling for confounding factors related to table size and answerability.
Appendix E Models
@@
Qwen2.5-7B 7.62B Sept 2024
Qwen2.5-7B-Instruct 7.62B Sept 2024
Qwen3-8B-Base 8.19B April 2025
Qwen3-8B 8.19B April 2025
Large Model(Larger than 2s0B parameter)
@@
Qwen2.5-72B 72.7B Sept 2024
Qwen2.5-72B-Instruct 72.7B Sept 2024
Qwen3-30B-A3B-Base 30.5B April 2025
Qwen3-30B-A3B 30.5B April 2025
@@
We selected recent open-source models that have been extensively studied and analyzed. Table 3 lists all the models we considered with their parameter size and their date of release. We include both base and instruction-tuned variants, allowing us to explore not only the effect of scale but also the impact of task specialization on tabular reasoning. The models span three major families: Llama-2, Llama-3, Mistral, Qwen2.5 and Qwen3, which together represent some of the most widely adopted transformer architectures in the open-source ecosystem. The inclusion of instruction-tuned variants is particularly important, as these models are optimized for following natural language instructions. This ability has been shown to influence performance on tasks that require a structured understanding significantly.
Moreover, by categorizing the models into ‘small’ and ‘large’ based on parameter count, we aim to systematically assess how model scale interacts with robustness and accuracy under various perturbation regimes. Recent models, such as Llama-3 and Mistral, demonstrate architectural innovations, including improved token representations and mixture-of-experts routing, which provide a richer set of inductive biases for our evaluation. Such a comprehensive suite enables a detailed comparison of robustness across architecture, scale, and fine-tuning strategies, thereby supporting more generalizable insights into LLM behavior on TQA tasks.
@@ Appendix F Performance over Table Complexity
Figure 13: The average EM score for the different structural perturbations over the different sizes of tables on WTQ, TAT-QA, and SCITAB datasets
@@
Across all three datasets, these results highlight that large language models, though capable, display uneven robustness to structural manipulations of tables. Although instruction tuning and larger model scales improve performance, structural changes continue to pose challenges, underscoring the need for more robust, structure-aware approaches to ensure reliable table comprehension.